# RILe: Reinforced Imitation Learning

### v1

@@ Abstract
Reinforcement Learning has achieved significant success in generating complex behavior but often requires extensive reward function engineering. Adversarial variants of Imitation Learning and Inverse Reinforcement Learning offer an alternative by learning policies from expert demonstrations via a discriminator. Employing discriminators increases their data- and computational efficiency over the standard approaches; however, results in sensitivity to imperfections in expert data. We propose RILe, a teacher-student system that achieves both robustness to imperfect data and efficiency. In RILe, the student learns an action policy while the teacher dynamically adjusts a reward function based on the student’s performance and its alignment with expert demonstrations. By tailoring the reward function to both performance of the student and expert similarity, our system reduces dependence on the discriminator and, hence, increases robustness against data imperfections. Experiments show that RILe outperforms existing methods by 2x in settings with limited or noisy expert data.
@@ 1 Introduction
Reinforcement Learning (RL) offers a framework for learning behavior by maximizing a reward function. In recent years, deep reinforcement learning has demonstrated remarkable success in replicating complex behaviors, including playing Atari games, chess, and Go [1, 2]. However, designing a reward function is a tedious and challenging task, as predicting the policy outcome from a manually crafted reward function is difficult.
To address this issue, Imitation Learning (IL) leverages expert demonstrations for learning a policy. Since vast amounts of expert data are required to learn expert behaviors accurately, Adversarial Imitation Learning (AIL) approaches are proposed as a data-efficient alternative to IL [3]. AIL employs a discriminator to measure similarity between learned behavior and expert behavior, and rewards the agent in training accordingly. However, the discriminator tends to overfit to the dynamics of the demonstration setting, limiting its ability to generalize to new environments [4]. Furthermore, AIL is vulnerable to noise or imperfections in expert data, as the discriminator rewards the agent to perfectly mimic expert data, including its flaws.
Inverse Reinforcement Learning (IRL) is another approach to alleviate reward engineering. In contrast to IL, which directly learns expert behavior, IRL seeks to infer the underlying reward function that leads to expert behavior. Thus, the reward function and the agent are trained iteratively, with updates to the reward function based on the agent’s behavior. This iterative process renders IRL computationally expensive [5]. Adversarial Inverse Reinforcement Learning (AIRL) [4] addresses this inefficiency by introducing a discriminator that allows for simultaneous learning of the policy and reward function. By implementing a special structure on the discriminator, AIRL can accurately recover the reward function of an expert, thereby improving generalization to environments with dynamics that vary from the demonstration setting, unlike AIL. However, in contrast to IRL, the reward function in AIRL is not learned independently but is directly derived from the output of the discriminator. Like AIL, AIRL suffers from the data-sensitivity issue associated with the discriminator, which limits its performance when expert data is imperfect.
To learn behaviors from imperfect data in a computationally efficient way, we propose Reinforced Imitation Learning (RILe). The primary motivation of RILe is to maintain computational efficiency and scalability through an adversarial framework, while learning a reward function from expert behavior (akin to IRL) as a separate mechanism. Our framework comprises two interacting agents: a student agent that learns a replicating policy and a teacher agent that learns a reward function and guides the student agent during training. For computational efficiency, we employ a discriminator that distinguishes expert data from student roll-outs as the reward function of the teacher. The core idea behind RILe is that the discriminator guides the reward learning process instead of directly guiding the action policy learning. Our contributions over AIL and AIRL are two-fold:
@@
    We learn the reward function (teacher agent) of the replicating policy (student agent) as an independent function in an adversarial IL/IRL setting. Using RL, the teacher agent learns to provide rewards that maximize the cumulative reward from the discriminator. The resulting reward function can provide meaningful guidance even when the discriminator overfits to noisy expert data, enabling the replicating policy to navigate and overcome sub-optimal states suggested by imperfect expert demonstrations.
@@
    The learned reward function in RILe adds an extra layer of flexibility to adversarial IL/IRL settings by allowing the teacher to customize rewards according to the developmental stage of the replicating agent, facilitating more accurate imitation of expert behavior (i.e., "online learned reward shaping").
Our experimental evaluation compares RILe to the state-of-the-art methods of AIL and AIRL, specifically GAIL [3] and AIRL [4], across three distinct sets of experiments: (1) imitating noisy expert data in a grid world, (2) imitating motion-capture data in continuous control tasks, and (3) learning to play image-based Atari games with discrete actions. Experimental results reveal that our approach outperforms baselines by 2x, especially when expert data is limited. We further demonstrate that RILe successfully learns expert-like behavior from noisy and misleading expert data, whereas the baseline methods fail to do so.
@@ 2 Related Work
We review the literature on learning expert behavior from demonstrations. Commonly, expert demonstrations are sourced either through direct queries to the expert in any observable state or by collecting sample trajectories demonstrated by the expert. We present related work that aligns with the most prevalent approaches of the latter setting, namely Imitation Learning and Inverse Reinforcement Learning. Both IL and IRL form the conceptual foundation of RILe.
Offline reinforcement learning also learns policies from data, which may include expert demonstrations. In contrast to our setting, its main goal is to learn a policy without any online interactions with the environment. We refer the reader to [6] for an overview of offline RL.
@@
The earliest work on imitation learning introduced Behavioral Cloning (BC) [7], which aims to learn a policy congruent with expert demonstrations through supervised learning. DAgger [8] proposes the aggregation of expert demonstrations with policy experiences during the training of the policy for improving generalization over expert demonstrations. GAIL introduces and adversarial imitation learning method, where a discriminator aims to understand whether queried behavior stems from a policy or from expert demonstrations, while a generator tries to fool the discriminator by learning a policy that exhibits expert-like behavior [3]. InfoGAIL extends upon GAIL by extracting latent factors from expert behavior and employing them during imitation learning [9]. Deep Q-learning from Demonstrations (DQfD) proposes to first pre-train the learning agent using expert demonstrations, followed by a subsequent policy optimization through interactions with the environment [10]. ValueDice introduces an off-policy imitation learning method using a distribution-matching objective between policy and expert behavior [11].
Although the field of imitation learning has seen advancements, the need for high-quality expert data and data efficacy remain open challenges [5]. Moreover, the limited generalization capability of IL approaches persists [12]. We address these limitations related to IL’s data efficacy and data sensitivity by introducing an intermediary teacher agent that adapts itself based on the current status of the student agent, enabling the teacher to effectively guide the student beyond the space covered by the expert state-action pairs.
@@
IRL is introduced in [13] to learn the intrinsic reward function of an expert and acquire the expert policy from it by iteratively optimizing an agents policy and then reward function. Apprenticeship learning builds on IRL and is proposed to represent the reward function as a linear combination of features [14]. Maximum Entropy Inverse Reinforcement Learning explicitly addresses the noise in expert demonstrations to better recover the reward function [15]. Several works extended IRL by incorporating negative demonstrations into the learning process to further improve behavior modeling [16, 17, 18]. Guided Cost Learning uses a neural network to approximate the reward function by using maximum entropy methods across continuous state-action spaces. [19]. AIRL proposes an adversarial reward learning framework to address the scalability issues of classical approaches [4]. IQ-Learn combines expert data with agent experiences to learn both reward and policy with a single Q-function [20]. Additionally, a pipeline is proposed that enables IRL to handle unstructured, real-world data [21]. XIRL introduces cross-embodiment scenarios, opening up a new direction in IRL research [22].
Despite the advancements in IRL, the computational efficacy of the learning process and scalability to complex problems remain open challenges [23]. The main reason for these limitations is the iterative sequential learning framework employed in IRL. We solve this efficacy problem by learning the policy and the reward function, via training a student agent and a teacher agent, in a single joint learning process.
@@ 3.2 Reinforcement Learning (RL)
Reinforcement learning seeks to find an optimal policy that maximizes the discounted cumulative reward given from the reward function $R=R(s,a)$. In this work, we incorporate entropy regularization using the $\gamma$-discounted casual entropy function $H(\pi)=E_{\pi}[-\text{log}\ \pi(a|s)]$ [3, 24]. The RL problem with a parameterized reward function and entropy regularization is defined as
@@ 3.3 Inverse Reinforcement Learning (IRL)
Given sample trajectories $\tau_{E}$ of an optimal expert policy $\pi_{E}$, inverse reinforcement learning $\text{IRL}(\tau_{E})$ seeks to recover a reward function $R_{\theta}^{*}(s,a)$ that maximally rewards behavior that was sampled from the expert policy $\pi_{E}$. In other words, IRL aims to find a reward function $R_{\theta}^{*}(s,a)$ that satisfies
$E_{\pi_{E}}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]\geq E_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]\quad\forall\pi.$ (2)
When IRL recovers the optimal reward function, $R_{\theta}^{*}(s,a)$ would result in a policy that exhibits expert behavior when optimized with reinforcement learning: $\text{RL}(R_{\theta}^{*}(s,a))=\pi^{*}=\operatorname*{argmax}_{\pi}E_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})].$
Since only the trajectories created by the expert policy are observed, expectations are estimated from the samples in $\tau_{E}$. In other words, IRL seeks to find the reward function in which the expert policy performs better than any other policy, where the expectation of the expert is estimated through samples. With entropy regularization $H(\pi)$, maximum casual entropy inverse reinforcement learning [15] can be defined as
@@
3.4 Adversarial Imitation Learning (AIL)
In contrast to inverse reinforcement learning, imitation learning aims to directly approximate the expert policy from given expert trajectory samples. It can be formulated as
@@ 3.4 Adversarial Imitation Learning (AIL)
GAIL [3] extends imitation learning to an adversarial setting by quantifying the difference between policies of the agent and the expert with a discriminator $D_{\phi}(s,a)$, parameterized by $\phi$. The discriminator’s function is to differentiate between expert-generated state-action pairs $(s,a)\sim\tau_{E}$ and non-expert state-action pairs $(s,a)\notin\tau_{E}$. Its optimization problem is defined as
$\max_{\phi}E_{(s,a)\sim\tau_{E}}[\text{log}(D_{\phi}(s,a))]+E_{(s,a)\notin\tau_{E}}[\text{log}(1-D_{\phi}(s,a))].$ (5)
Using the discriminator, the goal of GAIL is to find the optimal policy that minimizes this difference metric while maximizing an entropy constraint by training the discriminator and the policy at the same time. The optimization problem can be formulated as a zero-sum game between the discriminator $D_{\phi}(s,a)$ and the policy $\pi$, represented by
@@ 4 RILe: Reinforced Imitation Learning
We propose Reinforced Imitation Learning (RILe) to learn the reward function and acquire a policy that emulates expert-like behavior simultaneously in one learning process. Our framework consists of three key components: a student agent, a discriminator, and a teacher agent (Figure 1).
Intuitively, the student agent learns the action policy and the teacher agent learns the reward function. Both agents are trained simultaneously via RL with the help of an adversarial discriminator. The discriminator enables training the policy and reward function in one learning process and permits replicating expert behavior through similarity rather than supervised state action mapping. However, uniquely in RILe, the discriminator guides the reward learning process instead of directly guiding policy learning. This setup enables the teacher agent to learn a reward function separately by continuously adapting to changes in the student’s policy and guiding the student based on the discriminator’s feedback. As a result, the teacher can understand different regions of the state-action space through student’s interactions, effectively guiding it beyond the potentially noisy and imperfect expert state-action pairs.
@@
Figure 1: Framework overview. The framework consists of three key components: a student agent, a teacher agent, and a discriminator. (1) The student agent learns a policy $\pi_{S}$ by interacting with an environment. (2) Based on the state action pair of the student agent $s^{T}=(s^{S},a^{S})$, the teacher agent learns a reward function as a policy $\pi_{T}$ by (3) choosing an action $a^{T}$ that then becomes the reward of the student agent $a^{T}=r^{S}$. (4) The teacher agent forwards the $s^{T}=(s^{S},a^{S})$ to the discriminator. (5) The discriminator compares student state-action pair with expert demonstrations $s^{D}$ and (6) the rewards the teacher based on their similarity $r^{T}=D(s^{T})$.
@@
The student agent aims to learn a policy $\pi_{S}$ by interacting with an environment in a standard RL setting within an MDP. For each of its actions $a^{S}\in A$, the environment returns a new state $s^{S}\in S$. However, rather than from a hand-crafted reward function, the student agent receives its reward from the policy of the teacher agent $\pi_{T}$. Therefore, the reward function is represented by the teacher policy. Thus, the student agent is guided by the actions of the teacher agent, i.e., the action of the teacher is the reward of the student: $r^{S}=\pi_{T}((s^{S},a^{S}))$. The optimization problem of the student agent is then defined as
@@
Teacher Agent
The teacher agent aims to guide the student to approximate expert behavior by operating as its reward mechanism. Since the teacher agent takes the role of a reward function for the student, a new MDP is defined for the teacher agent: $\text{MDP}_{\text{T}}:(S_{T},A_{T},R_{T},T_{T},K,\gamma)$, where $S_{T}:S\ \text{x}\ A$ is the state space that consists all possible state action pairs from the standard MDP, defined in 3.1. $A_{T}$ is the action space, a mapping from $S_{T}\ \rightarrow\ \mathbb{R}$, so the action is a scalar value. $R_{T}(s,a)$ is the reward function where $s\in S_{T}$ and $a\in A_{T}$. $T_{T}=\{P(\cdot|s,a)\}$ is the transition dynamics where $P(\cdot|s,a)$ is defined as the state distribution upon taking action $a\in A_{T}$ in state $s\in S_{T}$. $K_{T}$ is the initial state distribution.
The teacher agent learns a policy $\pi_{T}$ that produces adequate reward signals to guide the student agent, by learning in a standard RL setting, within $MDP_{T}$. Since the state space of $MDP_{T}$ is defined over state-action pairs of $MDP_{S}$, the state of the teacher comprises the state-action pair of the student $s^{T}=(s^{S},a^{S})\in S_{T}$. It generates a scalar action $a^{T}$ which is given to the student agent as reward $r^{S}$, and bounded between -1 to 1. To help the teacher understand how its actions impact the reward it receives, we define the reward function such that it multiplies the discriminator’s output by a sigmoid function dependent on the teacher’s actions. Therefore, the teacher agent’s reward function is defined as $R^{T}=\text{log}(D_{\phi}(s^{T}))\sigma(a^{T})$, where $D_{\phi}(s^{T})$ is the output of the discriminator and $\sigma(a^{T})$ is the sigmoid function applied to the teacher’s own actions. Considering only the discriminator’s output in the teacher agent’s reward function would result in a reward independent of the teacher’s actions.
The optimization problem of the teacher can be defined as
$\max_{\pi_{T}}E_{\begin{subarray}{c}(s,a)\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}[\text{log}(D_{\phi}(s,a))\sigma(a^{T})].$ (9)
@@
RILe combines the three components defined previously in order to find a student policy that mimics expert behaviors presented in $\tau_{E}$. In RILe, the student policy $\pi_{S}$ and the teacher policy $\pi_{T}$ can be trained via any single-agent online reinforcement learning method. The training algorithm is given in Appendix C. Overall, the student agent aims to recover the optimal policy $\pi_{S}^{*}$ defined as

### v2

@@ Abstract
Reinforcement Learning has achieved significant success in generating complex behavior but often requires extensive reward function engineering. Adversarial variants of Imitation Learning and Inverse Reinforcement Learning offer an alternative by learning policies from expert demonstrations via a discriminator. However, these methods struggle in complex tasks where randomly sampling expert-like behaviors is challenging. This limitation stems from their reliance on policy-agnostic discriminators, which provide insufficient guidance for agent improvement, especially as task complexity increases and expert behavior becomes more distinct. We introduce RILe (Reinforced Imitation Learning environment), a novel trainer-student system that learns a dynamic reward function based on the student’s performance and alignment with expert demonstrations. In RILe, the student learns an action policy while the trainer, using reinforcement learning, continuously updates itself via the discriminator’s feedback to optimize the alignment between the student and the expert. The trainer optimizes for long-term cumulative rewards from the discriminator, enabling it to provide nuanced feedback that accounts for the complexity of the task and the student’s current capabilities. This approach allows for greater exploration of agent actions by providing graduated feedback rather than binary expert/non-expert classifications. By reducing dependence on policy-agnostic discriminators, RILe enables better performance in complex settings where traditional methods falter, outperforming existing methods by 2x in complex simulated robot-locomotion tasks.
@@ 1 Introduction
Reinforcement Learning (RL) has emerged as a powerful framework for teaching agents to perform complex tasks. In recent years, deep reinforcement learning has demonstrated remarkable success in replicating sophisticated behaviors, including playing Atari games, chess, and Go [1, 2]. However, these achievements often come at a cost: the tedious and challenging process of designing reward functions, as predicting the policy outcome from a manually crafted reward function remains notoriously difficult.
To overcome the reward engineering problem, Imitation Learning (IL) leverages expert demonstrations to learn a policy. Since vast amounts of expert data are required to accurately learn expert behaviors, Adversarial Imitation Learning (AIL) approaches, such as GAIL [3], have been proposed as data-efficient alternatives. AIL employs a discriminator to measure similarity between learned behavior and expert behavior, rewarding the agent accordingly. While computationally efficient, AIL methods suffer from a critical limitation: the policy-agnostic nature of their discriminators. The discriminator lacks any inherent incentive to guide the agent towards expert-like behavior, in contrast to engineered reward functions in RL. Consequently, AIL methods face challenges in complex tasks requiring extensive exploration to find optimal actions. For instance, in digital locomotion tasks, AIL methods often struggle to consistently replicate expert performance [4].
Inverse Reinforcement Learning (IRL) is another approach to alleviate reward engineering. Unlike IL, which directly learns expert behavior, IRL seeks to infer the underlying reward function that motivates the agent to acquire expert behaviors. The reward function and the agent are trained iteratively, with updates to the reward function based on the agent’s behavior. This iterative process renders IRL computationally expensive [5]. Adversarial Inverse Reinforcement Learning (AIRL) [6] attempts to address this inefficiency by introducing a discriminator that enables simultaneous learning of the policy and reward function. However, in AIRL, the reward function is tightly coupled to the discriminator, potentially limiting its ability to capture complex task structures or long-term dependencies and inheriting the limitations of a policy-agnostic discriminators. This highlights the need for a method that can learn a more flexible reward function without the computational overhead of traditional IRL methods.
To overcome these challenges and effectively learn behaviors in complex settings, we propose Reinforced Imitation Learning (RILe) (Fig. 1-(d)). RILe aims to combine the ability to learn a reward function that actively guides the agent to imitate expert behavior with the computational efficiency of adversarial frameworks. At the core of RILe is a novel trainer-student system designed to address the shortcomings of existing methods:
  • A student agent that learns to replicate the expert’s policy via RL in the environment
  • A trainer agent that learns a reward function via RL and guides the student agent during training
By integrating the trainer-student dynamic, RILe decouples reward learning from policy learning and the discriminator, allowing each component to specialize and thereby overcome the limitations of policy-agnostic discriminators. While RILe utilizes a discriminator similar to those in adversarial frameworks, its role is fundamentally redefined. In RILe, the discriminator’s primary function is to provide feedback to the trainer agent by distinguishing expert data from student roll-outs. This feedback serves as the reward signal for the trainer, not directly influencing the student agent. The trainer leverages the discriminator’s feedback to learn a reward function that effectively guides the student agent. This approach enables more nuanced reward shaping, particularly beneficial in tasks requiring complex decision-making and extensive exploration.
Our contributions are two-fold:
@@
    Decoupled Reward-function Learning: We introduce a novel approach where the trainer agent learns the reward function independently from both the student agent and the discriminator. Unlike existing methods that derive rewards directly from discriminator outputs, our trainer agent uses reinforcement learning to optimize the reward function based on the feedback from the discriminator. By focusing on long-term reward maximization, RL enables the trainer to distill inconsistent feedback from the discriminator into meaningful rewards, leading to better student performance.
@@
    Dynamic Reward Customization: Our trainer agent dynamically adjusts rewards based on the student agent’s progress, facilitating a better learning experience and enabling accurate imitation of expert behavior in complex settings. This adaptive approach allows for more gradual learning, particularly in tasks where the optimal behavior may change depending on the agent’s current capabilities.
We evaluate RILe against state-of-the-art methods in AIL, and AIRL, specifically GAIL [3], and AIRL [6]. Our experiments span three distinct scenarios: (1) Tailoring a reward function dynamically in a discrete maze task, (2) Investigating the impact of expert data on the trainer-student dynamics in a humanoid locomotion task, and (3) Imitating motion-capture data in continuous control tasks. The results demonstrate RILe’s superior performance, especially in complex tasks, and its ability to learn an effective dynamic reward function where baseline methods fail.
@@ 2 Related Work
We review literature on learning from expert demonstrations, focusing on Imitation Learning (IL) and Inverse Reinforcement Learning (IRL), which form the conceptual foundation of RILe.
@@
Early work introduced Behavioral Cloning (BC) [7], which learns a policy congruent with expert demonstrations through supervised learning. DAgger [8] introduces data aggregation. GAIL [3] introduces adversarial methods, where a discriminator aims to discriminate expert demonstrations, while a generator tries to fool the discriminator. DQfD [9] proposes two-stage approach with pre-training, and ValueDice [10] uses a distribution-matching objective between policy and expert. Despite progress, IL faces challenges in efficacy and generalization [5, 11]. RILe addresses these by introducing an adaptive teacher agent to guide the student beyond expert demonstrations.
@@
IRL, introduced by [12], learns the expert’s intrinsic reward function. Key developments include Apprenticeship Learning [13], Maximum Entropy IRL [14], and adversarial approaches like AIRL [6]. Recent work explores handling unstructured data [15] and cross-embodiment scenarios [16]. Despite advancements, IRL faces challenges in computational efficiency and scalability [17]. RILe addresses these by jointly learning policy and reward function in a single process.
@@ 3.2 Reinforcement Learning (RL)
Reinforcement learning seeks to find an optimal policy, $\pi^{*}$. that maximizes the discounted cumulative reward given from the reward function $R=R(s,a)$ (Fig. 1-(a)). In this work, we incorporate entropy regularization using the $\gamma$-discounted casual entropy function $H(\pi)=E_{\pi}[-\text{log}\ \pi(a|s)]$ [3, 19]. The RL problem with a parameterized reward function and entropy regularization is defined as
@@
(a) RL
(b) IRL
(c) GAIL + AIRL (terms in green)
(d) RILe
Figure 1: Overview of the related works. (a) Reinforcement Learning (RL): learning a policy that maximizes hand-defined reward function; (b) Inverse RL (IRL): learning a reward function from data. IRL has two stages: 1. training a policy with frozen reward function, and 2. updating the reward function by comparing the converged policy with data. These stages repeated several times; (C) Generative Adversarial Imitation Learning (GAIL) + Adversarial IRL (AIRL): using discriminator as a reward function. GAIL trains both policy and the discriminator at the same time. AIRL implements a new structure on the discriminator, seperating reward from environment dynamics by using two networks under the discriminator (see additional terms in green). (D) RILe: similar to IRL, learning a reward function from data. RILe learns the reward function at the same time with the policy, using discriminator as a guide for learning the reward.
@@ 3.3 Inverse Reinforcement Learning (IRL)
Given sample trajectories $\tau_{E}$ from an optimal expert policy $\pi_{E}$, inverse reinforcement learning aims to recover a reward function $R^{*}_{\theta}(s,a)$ that maximally rewards the expert’s behavior (Fig. 1-(b)). Formally, IRL seeks a reward function, $R_{\theta}^{*}(s,a)$, satisfying: $E_{\pi_{E}}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]\geq E_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]\quad\forall\pi.$ Optimizing this reward function with reinforcement learning yields a policy that replicates expert behavior: $\text{RL}(R_{\theta}^{*}(s,a))=\pi^{*}=\operatorname*{arg\,max}_{\pi}E_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]$.
Since only the expert’s trajectories are observed, expectations over $\pi_{E}$ are estimated from samples in $\tau_{E}$. Incorporating entropy regularization $H(\pi)$, maximum causal entropy inverse reinforcement learning [14] is defined as
@@
3.4 Adversarial Imitation Learning (AIL) and Adversarial Inverse Reinforcement Learning (AIRL)
Imitation Learning (IL) aims to directly approximate the expert policy from given expert trajectory samples $\tau_{E}$. It can be formulated as
@@ 3.4 Adversarial Imitation Learning (AIL) and Adversarial Inverse Reinforcement Learning (AIRL)
GAIL [3] introduces an adversarial imitation learning setting by quantifying the difference between the agent and the expert with a discriminator $D_{\phi}(s,a)$, parameterized by $\phi$ (Fig. 1-(c)).
The discriminator distinguishes between between expert-generated state-action pairs $(s,a)\sim\tau_{E}$ and non-expert state-action pairs $(s,a)\notin\tau_{E}$. The goal of GAIL is to find the optimal policy that fools the discriminator while maximizing an entropy constraint. The optimization is formulated as a zero-sum game between the discriminator $D_{\phi}(s,a)$ and the policy $\pi$:
@@
AIRL [6] extends this adversarial approach to inverse reinforcement learning, aiming to recover a reward function decoupled from environment dynamics (Fig. 1-(c)). The key insight of AIRL is to structure the discriminator as:
$D_{\phi,\psi}(s,a,s^{\prime})=\frac{\exp(f_{\phi}(s,a,s^{\prime}))}{\exp(f_{\phi}(s,a,s^{\prime}))+\pi(a|s)},$ (5)
where $f_{\phi}(s,a,s^{\prime})=r_{\psi}(s,a)+\gamma V_{\phi}(s^{\prime})-V_{\phi}(s)$. Here, $r_{\psi}(s,a)$ represents the learned reward function that is decoupled from the environment dynamics, $\gamma V_{\phi}(s^{\prime})-V_{\phi}(s)$. The AIRL optimization problem is formulated equivalently to GAIL (see Eqn. 4). The reward function $r_{\psi}(s,a)$ is learned through minimizing the cross-entropy loss inherent in this adversarial setup. Therefore, the reward function remains tightly coupled with the discriminator’s learning process.
@@ 4 RILe: Reinforced Imitation Learning
We propose Reinforced Imitation Learning (RILe) to learn the reward function and acquire a policy that emulates expert-like behavior simultaneously in one learning process. Our framework consists of three key components: a student agent, a trainer agent, and a discriminator (Fig. 2).
Intuitively, the student agent learns the action policy and the trainer agent learns the reward function. Both agents are trained simultaneously via RL with the help of an adversarial discriminator. Building on AIL, we utilize the discriminator to enable joint training of policy and reward function. Unlike traditional IRL, the discriminator provides immediate feedback to the trainer, eliminating the need to wait for the student agent’s convergence.
The trainer agent, trained via RL, is rewarded by the discriminator, and thus motivated to maximize the cumulative reward from the discriminator. The trainer learns to craft a reward function that promotes expert-like behavior in the student, even in complex scenarios. Crucially, the trainer’s RL-based approach optimizes for long-term performance rather than immediate feedback. This allows the trainer to encourage exploration and gradual improvement in three main ways. First, the trainer can associate its reward-giving actions with future improvements in the student’s behavior, even if those improvements take many steps to manifest. Second, the trainer can provide small positive rewards for actions that move the student in the right direction, even if they don’t immediately result in expert-like behavior. Finally, the trainer can adapt its reward function based on the student’s current capabilities, providing a learning path that gradually guides the student towards expert behavior. This approach enables RILe to overcome limitations of previous methods, particularly in complex tasks requiring extensive exploration, by promoting the discovery of expert-like strategies even when the student’s initial policy significantly diverges from expert behavior.
@@
Figure 2: Framework overview. The framework consists of three key components: a student agent, a trainer agent, and a discriminator. The student agent learns a policy $\pi_{S}$ by interacting with an environment, and the trainer agent learns a reward function as a policy $\pi_{T}$. (1) The student receives the environment state $s^{S}$. (2) The student takes an action $a^{S}$, forwards it to the environment which is updated based on $a^{S}$. (3) The student forwards its state and action to the trainer, whose state is $s^{T}=(s^{S},a^{S})$. (4) Trainer, $\pi_{T}$, evaluates the state action pair of the student agent $s^{T}=(s^{S},a^{S})$ and chooses an action $a^{T}$ that then becomes the reward of the student agent $a^{T}=r^{S}$. (5) The trainer agent forwards the $s^{T}=(s^{S},a^{S})$ to the discriminator. (6) Discriminator compares student state-action pair with expert demonstrations ($s^{D}$). (7) Discriminator gives reward to the trainer, based on the similarity between student- and expert-behavior.
@@
The student agent aims to learn a policy $\pi_{S}$ by interacting with an environment in a standard RL setting within an MDP. For each of its actions $a^{S}\in A$, the environment returns a new state $s^{S}\in S$. However, rather than from a hand-crafted reward function, the student agent receives its reward from the policy of the trainer agent $\pi_{T}$. Therefore, the reward function is represented by the trainer policy. Thus, the student agent is guided by the actions of the trainer agent, i.e., the action of the trainer is the reward of the student: $r^{S}=\pi_{T}((s^{S},a^{S}))$. The optimization problem of the student agent is then defined as
@@
Trainer Agent
The trainer agent aims to guide the student to approximate expert behavior by operating as its reward mechanism. Since the trainer agent takes the role of a reward function for the student, a new MDP is defined for the trainer agent: $\text{MDP}_{\text{T}}:(S_{T},A_{T},R_{T},T_{T},K,\gamma)$, where $S_{T}:S\ \text{x}\ A$ is the state space that consists all possible state action pairs from the standard MDP, defined in 3.1. $A_{T}$ is the action space, a mapping from $S_{T}\ \rightarrow\ \mathbb{R}$, so the action is a scalar value. $R_{T}(s,a)$ is the reward function where $s\in S_{T}$ and $a\in A_{T}$. $T_{T}=\{P(\cdot|s,a)\}$ is the transition dynamics where $P(\cdot|s,a)$ is defined as the state distribution upon taking action $a\in A_{T}$ in state $s\in S_{T}$. $K_{T}$ is the initial state distribution.
The trainer agent learns a policy $\pi_{T}$ that produces adequate reward signals to guide the student agent, by learning in a standard RL setting, within $MDP_{T}$. Since the state space of $MDP_{T}$ is defined over state-action pairs of $MDP_{S}$, the state of the trainer comprises the state-action pair of the student $s^{T}=(s^{S},a^{S})\in S_{T}$. It generates a scalar action $a^{T}$ which is given to the student agent as reward $r^{S}$, and bounded between -1 to 1. To help the trainer understand how its actions impact the reward it receives, we define the reward function such that it multiplies the scaled discriminator’s output by trainer’s actions. Therefore, the trainer agent’s reward function is defined as $R^{T}=\upsilon(D_{\phi}(s^{T}))(a^{T})$, where $D_{\phi}(s^{T})$ is the output of the discriminator and $\upsilon(x)=2x-1$ is the scaling function. By incorporating $a^{T}$ into the reward function, the trainer learns to adjust its policy based on the effectiveness of its previous actions.
The optimization problem of the trainer can be defined as
$\max_{\pi_{T}}E_{\begin{subarray}{c}(s,a)\sim\pi_{S}\\ a^{T}\sim\pi_{T}\end{subarray}}[\upsilon(D_{\phi}(s,a))a^{T}].$ (8)
@@
RILe combines the three components defined previously in order to find a student policy that mimics expert behaviors presented in $\tau_{E}$. In RILe, the student policy $\pi_{S}$ and the trainer policy $\pi_{T}$ can be trained via any single-agent online reinforcement learning method. The training algorithm is given in Appendix D. Overall, the student agent aims to recover the optimal policy $\pi_{S}^{*}$ defined as